OpenStudy (anonymous):

Which shows 542 - 462 being evaluated using the difference of squares method?

OpenStudy (legends):

Is there any answers

OpenStudy (anonymous):

Which shows 542 - 462 being evaluated using the difference of squares method? 542 - 462 = (2916 + 2116)(2916 - 2116) = 4,025,600 542 - 462 = (54 + 46)(54 - 46) = (100)(8) = 800 542 - 462 = 2916 - 2116 = 800 542 - 462 = (54 - 46)2 = 82 = 64 sorry

OpenStudy (anonymous):

what do you think @blushy18

OpenStudy (legends):

Im not 100% sure but I think its the second one idk for sure tho.

OpenStudy (legends):

Or the third one idk

OpenStudy (legends):

@blushy18 You there?